# Tunable electron-electron interaction and anomalous enhancement of large-momentum scattering in moiré superlattices

### v3

@@ Abstract
Using a microscopic model, we show that the electron-electron interaction of flat bands deviates significantly from the Coulomb interaction. In particular, we find that large-momentum scattering is enhanced at $\theta\lesssim 4^{\circ}$, with a non-monotonic momentum dependence appearing near the magic angle. For $\theta\gtrsim 1.2^{\circ}$, the enhanced large-momentum scattering can be attributed to the compact Wannier function. On the other hand, for $\theta\lesssim 1.2^{\circ}$, the nonmonotonic momentum dependence of the interaction matrix cannot be explained by a simple Wannier orbital, indicating a nontrivial modification to the el-el interaction. Notably, the range of angles $\theta$ where the large-momentum scattering is enhanced differs from the magic angles at which nearly-flat bands emerge, suggesting that the angle dependence of material properties provides information about the effect of interaction. The results highlight unusual features of the interaction in moiré graphene.
@@ I Introduction
In this work, we investigate the el-el interaction in moiré graphene using a microscopic theory based on a tight-binding model [7, 39]. We show that, beyond its magnitude, the structure of the el-el interaction matrix for the flat bands undergoes significant changes with the twist angle [Fig. 1(c)]. Specifically, we find that large-momentum scattering is enhanced compared to the bare Coulomb interaction. The enhancement of large-momentum scattering at twist angles $\theta\lesssim 1.2^{\circ}$ shows a nonmonotonic $q$ dependence, which is qualitatively different from the typical cases. For the cases $\theta\gtrsim 1.3^{\circ}$, the interaction matrix has a form qualitatively similar to the usual Coulomb interaction. However, a considerable increase of the large-$q$ scattering occurs in a wide range of twist angles up to a few degrees. Our results indicate that, in addition to the emergence of narrow electronic bands, moiré graphene hosts a highly nontrivial el-el interaction structure that is strongly influenced by the underlying Wannier orbitals, with potential implications for its correlated phases.
@@ II.2 Interaction matrix
To investigate the nature of the el-el interaction, we calculated the interaction matrix for the flat bands. To this end, we consider the Yukawa interaction between the electrons,
@@
is the interaction matrix. In Eq. (5), the sum is taken over all momenta $\mathbf{q}^{\prime}$ which are equivalent to $\mathbf{q}$ in the first moiré Brillouin zone.
@@
Figure 2: The band and interaction matrix of electron-electron interaction. (a) The dispersion $E_{\mathbf{k}}$ of flat bands $n=1,2,3,4$ for twist angle $\theta=1.16^{\circ}$ calculated along the path $B$ (see inset of (b)). The even and odd bands belong to different valleys. The interaction matrix for twist angles (b) $\theta=3.89^{\circ}$, (c) $1.47^{\circ}$, and (d) $1.05^{\circ}$. Inset of (b) shows three paths along which the interaction matrix was calculated: (A) $\mathbf{k}=29/30\mathbf{K}_{1}+1/30\mathbf{K}_{2}$ and $\mathbf{k}^{\prime}=1/30\mathbf{K}_{1}+29/30\mathbf{K}_{2}$, (B) $\mathbf{k}=29/30\mathbf{K}_{1}-28/30\mathbf{K}_{2}$ and $\mathbf{k}^{\prime}=1/30\mathbf{K}_{1}+28/30\mathbf{K}_{2}$, and (C) $\mathbf{k}=29/30\mathbf{K}_{1}-28/30\mathbf{K}_{2}$ and $\mathbf{k}^{\prime}=-28/30\mathbf{K}_{1}+29/30\mathbf{K}_{2}$. In all cases, the direction of $\mathbf{q}$ is taken so that $\overline{\mathbf{k-q}}$ moves from $\mathbf{k}$ to $\mathbf{k^{\prime}}$.
@@ III Results
On the other hand, $V_{\mathbf{k},\mathbf{k}^{\prime},\mathbf{q}}^{n,n^{\prime},m,m^{\prime}}$ along paths $A$ and $C$ show nonmonotonic $q$ dependence near the magic angle, as seen in the results for $\theta=1.05^{\circ}$ in Fig. 2(d). The results show that the large angular momentum by the el-el interaction is enhanced near the magic angle, indicating a unique feature of the el-el interaction in moiré graphene. The enhancement of large momentum scattering is consistent with what is predicted in a previous work [37]. However, the nonmonotonic $q$ dependence was not predicted. As shown in Fig. 1(c), the enhanced large-momentum scattering is seen in a range of angles around $\theta=1.1^{\circ}$.
@@ IV Conclusion
In this work, we studied the electron-electron interaction matrix using a transferrable tight-binding model for carbon atoms and by constructing an effective tight-binding model. Using the transferrable tight-binding model, we find that the large-$q$ scattering by electron-electron interaction is enhanced in the flat bands. For the twist angles $\theta\gtrsim 1.2$, the enhancement of large-$q$ scattering is well reproduced by an effective compression of $\xi_{G}$; $\xi_{G}$ is smaller than the unit cell, reaching below $1/4$ of the length of primitive vectors at $\theta<2^{\circ}$. The enhancement of large-$q$ scattering by a Wannier orbital with a small $\xi_{G}$ was predicted in a recent work [37]. On the other hand, for $\theta\lesssim 1.2^{\circ}$ we find a non-monotonic $q$ dependence of the interaction matrix. Such a $q$ dependence cannot be explained by the effective tight-binding model. As the non-monotonic $q$ dependence of electron-electron interaction is seen only at small twist angles near the first magic angle, it is possibly related to the unusual electronic properties of moiré graphene. The enhanced electron-electron interaction taking place at $\theta\lesssim 2^{\circ}$, which is tunable by controlling the twist angle may provide another way of controlling the correlation effect.
@@
The enhancement of large-$q$ scattering indicates that careful consideration of the interaction matrix might be necessary to understand the interaction physics of moiré graphene, in addition to the narrow bandwidth. The enhanced large-$q$ scattering is likely to affect the transport properties by enhancing the scattering of electrons by the electron-electron interaction. Interestingly, such enhancement of the electron-electron interaction is seen below $\theta\sim 4^{\circ}$. This is different from the flat band formation, which is highly sensitive to the twist angle; the bandwidth is more than ten times larger than that of the magic angle at $\theta\sim 1.1^{\circ}$. Therefore, the twist-angle dependence of the material properties can be used to investigate the origin of the interaction physics in moiré graphene.
@@
On the other hand, the unusual $q$ dependence of electron-electron interaction near the magic angle was not reproducible within the phenomenological model considered. One possible explanation for this could be the clover-leaf-like structure of the Wannier orbital [45]. The form of the Wannier function significantly affects the interaction matrix in momentum space. The clover-leaf-like structure of the Wannier orbital may contribute to the unusual $q$ dependence. Another cause of failure might be the Wannier obstruction [47, 48, 49, 50]. The Wannier obstruction prohibits constructing exponentially decaying Wannier orbitals. As the $q$ dependence of the interaction matrix is affected by the functional form of the Wannier orbitals, the form of $V_{\mathbf{k},\mathbf{k}^{\prime},\mathbf{q}}^{n,n^{\prime},m,m^{\prime}}$ may be significantly affected by the Wannier obstruction. If such an effect exists, assuming the interaction matrix to be similar to that of the model in Eq. 8 might be unsuitable. Therefore, a model incorporating the Wannier obstruction, such as those in Refs. [28, 44] might provide a better starting point. Regardless of the origin, the unusual $q$ dependence is likely to affect the low-temperature property of moiré graphene, through enhanced large $q$ scattering. The results and analysis presented in this work unveil nontrivial modifications of the electron-electron interaction, which affect the correlation physics in moiré graphene.

### v4

# Anomalous enhancement of large-momentum scattering by electron-electron interaction in moiré superlattices

@@ Abstract
Using a microscopic model, we show that the large-momentum scattering of electrons in flat bands deviates significantly from that of the Coulomb scattering. In particular, we find that large-momentum scattering is enhanced at $\theta\lesssim 4^{\circ}$, with a non-monotonic momentum dependence appearing near the magic angle. For $\theta\gtrsim 1.2^{\circ}$, the enhanced large-momentum scattering can be attributed to the compact Wannier function. On the other hand, for $\theta\lesssim 1.2^{\circ}$, the nonmonotonic momentum dependence of the interaction matrix cannot be explained by a simple Wannier orbital, indicating a nontrivial modification of the el-el scattering. Notably, the range of angles $\theta$ where the large-momentum scattering is enhanced differs from the magic angles at which nearly-flat bands emerge, suggesting that the angle dependence of material properties provides information about the effect of interaction. The results highlight unusual features of the interaction in moiré graphene.
@@ I Introduction
In this work, we investigate the el-el interaction matrix in moiré graphene using a microscopic theory based on a tight-binding model [7, 39]. We show that, beyond its magnitude, the structure of the el-el interaction matrix for the flat bands undergoes significant changes with the twist angle [Fig. 1(c)]. Specifically, we find that large-momentum scattering is enhanced compared to the bare Coulomb interaction. The enhancement of large-momentum scattering at twist angles $\theta\lesssim 1.2^{\circ}$ shows a nonmonotonic $q$ dependence, which is qualitatively different from the typical cases. For the cases $\theta\gtrsim 1.3^{\circ}$, the interaction matrix has a form qualitatively similar to the usual Coulomb interaction. However, a considerable increase of the large-$q$ scattering occurs in a wide range of twist angles up to a few degrees. Our results indicate that, in addition to the emergence of narrow electronic bands, moiré graphene hosts a highly nontrivial el-el scattering that is strongly influenced by the underlying Wannier orbitals, with potential implications for its correlated phases.
@@ II.2 Interaction matrix
To investigate the nature of the el-el scattering, we calculated the interaction matrix for the flat bands. To this end, we consider the Yukawa interaction between the electrons,
@@
is the interaction matrix. The interaction matrix corresponds to electron-electron scattering. In Eq. (5), the sum is taken over all momenta $\mathbf{q}^{\prime}$ which are equivalent to $\mathbf{q}$ in the first moiré Brillouin zone.
@@
Figure 2: The band and interaction matrix of flat-band electrons. (a) The dispersion $E_{\mathbf{k}}$ of flat bands $n=1,2,3,4$ for twist angle $\theta=1.16^{\circ}$ calculated along the path $B$ (see inset of (b)). The even and odd bands belong to different valleys. The interaction matrix for twist angles (b) $\theta=3.89^{\circ}$, (c) $1.47^{\circ}$, and (d) $1.05^{\circ}$. Inset of (b) shows three paths along which the interaction matrix was calculated: (A) $\mathbf{k}=29/30\mathbf{K}_{1}+1/30\mathbf{K}_{2}$ and $\mathbf{k}^{\prime}=1/30\mathbf{K}_{1}+29/30\mathbf{K}_{2}$, (B) $\mathbf{k}=29/30\mathbf{K}_{1}-28/30\mathbf{K}_{2}$ and $\mathbf{k}^{\prime}=1/30\mathbf{K}_{1}+28/30\mathbf{K}_{2}$, and (C) $\mathbf{k}=29/30\mathbf{K}_{1}-28/30\mathbf{K}_{2}$ and $\mathbf{k}^{\prime}=-28/30\mathbf{K}_{1}+29/30\mathbf{K}_{2}$. In all cases, the direction of $\mathbf{q}$ is taken so that $\overline{\mathbf{k-q}}$ moves from $\mathbf{k}$ to $\mathbf{k^{\prime}}$.
@@ III Results
On the other hand, $V_{\mathbf{k},\mathbf{k}^{\prime},\mathbf{q}}^{n,n^{\prime},m,m^{\prime}}$ along paths $A$ and $C$ show nonmonotonic $q$ dependence near the magic angle, as seen in the results for $\theta=1.05^{\circ}$ in Fig. 2(d). The results show that the large angular momentum scattering by the el-el interaction is enhanced near the magic angle, indicating a unique feature of the el-el scattering in moiré graphene. The enhancement of large momentum scattering is consistent with what is predicted in a previous work [37]. However, the nonmonotonic $q$ dependence was not predicted. As shown in Fig. 1(c), the enhanced large-momentum scattering is seen in a range of angles around $\theta=1.1^{\circ}$.
@@ IV Conclusion
In this work, we studied the electron-electron interaction matrix using a transferrable tight-binding model for carbon atoms and by constructing an effective tight-binding model. Using the transferrable tight-binding model, we find that the large-$q$ scattering by electron-electron interaction is enhanced in the flat bands. For the twist angles $\theta\gtrsim 1.2$, the enhancement of large-$q$ scattering is well reproduced by an effective compression of $\xi_{G}$; $\xi_{G}$ is smaller than the unit cell, reaching below $1/4$ of the length of primitive vectors at $\theta<2^{\circ}$. The enhancement of large-$q$ scattering by a Wannier orbital with a small $\xi_{G}$ was predicted in a recent work [37]. On the other hand, for $\theta\lesssim 1.2^{\circ}$ we find a non-monotonic $q$ dependence of the interaction matrix. Such a $q$ dependence cannot be explained by the effective tight-binding model. As the non-monotonic $q$ dependence of the interaction matrix is seen only at small twist angles near the first magic angle, it is possibly related to the unusual electronic properties of moiré graphene. The enhanced electron-electron scattering taking place at $\theta\lesssim 2^{\circ}$, which is tunable by controlling the twist angle may provide another way of controlling the correlation effect.
@@
The enhancement of large-$q$ scattering indicates that careful consideration of the interaction matrix might be necessary to understand the interaction physics of moiré graphene, in addition to the narrow bandwidth. Considering that the electrons near the Fermi surface dominantly contribute to the transport, the enhanced large-$q$ scattering is likely to affect the transport properties of the moiré graphene. Interestingly, such enhancement of the electron-electron scattering is seen below $\theta\sim 4^{\circ}$. This is different from the flat band formation, which is highly sensitive to the twist angle; the bandwidth is more than ten times larger than that of the magic angle at $\theta\sim 1.1^{\circ}$. Therefore, the twist-angle dependence of the material properties can be used to investigate the origin of the interaction physics in moiré graphene.
@@
On the other hand, the unusual $q$ dependence of the interaction matrix near the magic angle was not reproducible within the phenomenological model considered. One possible explanation for this could be the clover-leaf-like structure of the Wannier orbital [45]. The form of the Wannier function significantly affects the interaction matrix in momentum space. The clover-leaf-like structure of the Wannier orbital may contribute to the unusual $q$ dependence. Another cause of failure might be the Wannier obstruction [47, 48, 49, 50]. The Wannier obstruction prohibits constructing exponentially decaying Wannier orbitals. As the $q$ dependence of the interaction matrix is affected by the functional form of the Wannier orbitals, the form of $V_{\mathbf{k},\mathbf{k}^{\prime},\mathbf{q}}^{n,n^{\prime},m,m^{\prime}}$ may be significantly affected by the Wannier obstruction. If such an effect exists, assuming the interaction matrix to be similar to that of the model in Eq. 8 might be unsuitable. Therefore, a model incorporating the Wannier obstruction, such as those in Refs. [28, 44] might provide a better starting point. Regardless of the origin, the unusual $q$ dependence is likely to affect the low-temperature property of moiré graphene, through enhanced large $q$ scattering. The results and analysis presented in this work unveil nontrivial modifications of the electron-electron scattering, which affect the correlation physics in moiré graphene.